OpenStudy (anonymous):

Need some help please: How many pounds of cashews selling at 4$ per pound should be mixed with 3 pounds of peanuts selling at $1.60 per pound to make a mixture selling for $2.80 per pound?

OpenStudy (stamp):

Identify: c = pounds of cashews p = pounds of peanuts p = 3 m = pounds of peanuts and cashews mixed Setup: 4c + 1.60p = 2.80m c + p = m Solve

OpenStudy (anonymous):

I am not seeing where this makes much sence haha brain is kind of fried.

OpenStudy (stamp):

4c + 1.60p = 2.80m c + p = m --- p = 3 4c + 1.60(3) = 2.80m c + (3) = m ------------- c = m - 3 4(m-3) + 1.60(3) = 2.80m 4m - 12 + 4.80 = 2.80m 4m - 7.20 = 2.80 m 1.20m - 7.20 = 0 1.20m = 7.20 m = 7.20 / 1.20 m = 6 --- m = pounds of mixture = 6 lbs recall: c + p = m c + 3 = 6 c = 3 lbs

OpenStudy (stamp):

It is algebra, you assign variables to your elements and use your math skills and tricks to solve for the variables. Substitution, distributive property, etc. If you are not comfortable or familiar with these definitions, you absolute have to go back and study. Look them up online and do some practice, go back into your old textbooks and practice. And once you have practiced them some, keep practicing.

OpenStudy (anonymous):

I understand whereI was getting confused thanks for your help.